Shift report 20030502 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri May 02 2003 - 15:35:11 EDT

  • Next message: Apache: "Shift report 20030502 16:00-24:00"
    ------------------------------------------------------------
    Date: Fri May  2 15:35:10 2003, Shift: 08:00-16:00
    Supervisor: EUn-Joo Kim, with: 
    Log book#14 pages 90 - 90
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam for most shift time.
       RF setup work continued.
       Spin pattern is not complete and so on.
    
    2. FS 1/8A at 8-degree
       -- Run8764, 8765 : 39k trigger2
       FS D3,D4 zero-field at 8-deg
       -- Run8766 : 11k trigger2
       FS 1/8B at 8-degree
       -- Run8768, 8769 : 13k trigger2
    
       *** Run8767 : special run for FV  
            The spec triggers are the hodoscopes + RC . 
            The INL was excluded.
            The INL trigger ,i.e.,5 is L.R (w/o) RC.
    
       MRS 6B at 40-degree
        -- Run8764, 8765, 8766: 245K trigger3
       MRS 6A at 40-degree
        -- Run8768, 8769 : 57k trigger3
    
    3. Mainframe for TFW2 : still not working
       
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Fri May 02 2003 - 15:35:51 EDT